OpenStudy (anonymous):

Please help!!!! Fan and medal.Use the inverse properties of logarithms to simplify the expression.10^(log21)

jimthompson5910 (jim_thompson5910):

Use the same idea as last time, just with a different base this time (10 instead of e)

jimthompson5910 (jim_thompson5910):

Rule: \[\LARGE 10^{\log(x)} = x\] \[\LARGE \log\left(10^x\right) = x\]

OpenStudy (anonymous):

So replace the x with 21? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

then the answer is 21?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep 21

OpenStudy (anonymous):

thank you very much

jimthompson5910 (jim_thompson5910):

yw